# Fire [Uncaptioned image]: Fact-checking with Iterative Retrieval and Verification

### v1

@@ 1 Introduction
Large language models (LLMs) have demonstrated exceptional performance across a wide range of tasks, including both language comprehension and generation (Zhao et al., 2023; Xie et al., 2023a). Consequently, LLMs are now widely applied in various domains (Xie et al., 2023b), and many users increasingly rely on the information they provide. However, this reliance is problematic, as LLMs are capable of producing outputs that are highly confident but factually incorrect, highlighting the critical need for robust fact-checking systems (Akhtar et al., 2023). However, fact-checking the entire output of LLMs in a single step is highly challenging. To address this, Min et al. (2023) proposed decomposing the content into multiple atomic claims, each of which can be individually verified. While this approach simplifies the fact-checking process, assessing the veracity of these atomic claims remains complex, especially when many require sourcing evidence from the web. Indeed, identifying the most relevant evidence online is a key challenge in fact-checking pipelines (Wang et al., 2023).
To address this issue, conventional methods, such as FacTool and Factcheck-GPT (Chern et al., 2023; Wang et al., 2023), frame the problem as a question-answering task, as illustrated on the left side of Figure 1. In these approaches, an LLM is prompted to generate N relevant questions, which are then used as search queries by a web search tool. The search results serve as evidence for LLM to determine the factuality of the claim. However, we argue that this process is inefficient in two key aspects. First, it underutilizes the internal knowledge already embedded in LLMs during pre-training. For claims involving common knowledge or widely known events, the LLM could confidently assess the claim without relying on external information. Second, generating multiple search queries concurrently does not align with the typical human reasoning process during search (Hu et al., 2023). Humans tend to begin with an initial query, gather information, and then refine their perspective on the claim, which often leads to the formulation of more effective follow-up queries.
@@ 2 Related Work
Despite the remarkable capabilities of LLMs (Brown et al., 2020; OpenAI, 2023; Zhao et al., 2023), the auto-regressive learning objective does not inherently offer strong guarantee or enforce the learning of factual accuracy in the training process, making these models produce content that deviates from real-world facts (Wang et al., 2024a). On average, there are 5%-10% false claims in responses of GPT-4 (OpenAI, 2023) and LLaMA-2 (Touvron et al., 2023) on world-knowledge questions (Iqbal et al., 2024). Retrieval-augmented generation (Guu et al., 2020) and post-generation fact-checking are essential for ensuring accurate knowledge dissemination. Retrieving highly relevant information plays a pivotal role in both guiding generation as a reference and determining verification results in fact-checking systems (Wang et al., 2023).
@@
The recent advancements in LLMs have spurred significant research on LLM-powered agents, which are capable of reasoning about their environment and making decisions by either invoking external tools or performing internal actions (Wang et al., 2024b). These agent frameworks typically consist of several components, including reasoning, tool usage, memory, and multi-agent debate (Masterman et al., 2024), many of which can be seamlessly integrated into fact-checking pipelines to enhance the performance of traditional fact-checking systems. For example, recent works have endowed systems with the ability to call external tools, such as search engines (Chern et al., 2023; Wang et al., 2023; Wei et al., 2024; Cheng et al., 2024), recognizing that many claims in the field require additional information for verification. During the verification stage, Sun et al. (2024) proposed a Markov Chain-based multi-agent debate approach to ensure more rigorous verification by enabling collaborative decision-making among agents based on retrieved evidence. Our work differs from previous approaches by combining the evidence retrieval and verification stages, leveraging agents’ reasoning and tool-use capabilities to more closely simulate human cognitive processes in fact-checking.
@@ 3 Framework
Assessing the factual accuracy of long-form text presents significant challenges (Min et al., 2023). To address this, prior approaches have broken down the text into individual checkworthy claims (Chern et al., 2023). These sentences, referred to as atomic claims, are fact-checked individually, with their factuality scores aggregated to evaluate the overall factual accuracy of the original text. Previous research indicates that verifying the factuality of atomic claims is the most challenging step in this process (Wang et al., 2023). Our work therefore focuses on this critical task: determining the factual accuracy of individual atomic claims, classifying each as either True or False.
@@ 3.1 Fire
We introduce a unified method, Final Answer or Next Search Query $f(\cdot)$, which integrates claim verification with search query generation. This component processes an atomic claim $c$ and determines whether to produce a final answer $a$ or to generate an additional search query $q$. This decision is informed by both an external evidence set $E$, derived from search results, and the internal knowledge $k$ of the language model, acquired during pre-training. To facilitate this, we incorporate confidence estimation into the reasoning process. The model outputs a final answer $a$ when its confidence is sufficiently high; otherwise, it generates an additional query $q$. This process can be formalized as follows:
$f(c,E,k)=\begin{cases}a,&\text{ if confident}\\ q,&\text{ if not confident}\end{cases}$
@@
When the language model determines that a web search is necessary by issuing a search query $q$, we utilize Google Search through the SerpAPI [^1] https://serpapi.com. This service returns the search results as new evidence $e$, which we then append to the existing evidence set $E$ to form the new evidence set $E^{\prime}$ for the next iteration.
$E^{\prime}=E\cup e,e=\text{Search}(q)$
@@
Due to the inherent difficulty in confidently verifying certain claims, even after retrieving supplementary evidence, we establish an upper limit on the number of retrieval steps. Upon reaching this limit, the model is required to conduct a final verification $f^{\prime}(\cdot)$ based on all previously retrieved evidence. The detailed prompt for this process is provided in Appendix B.
$\begin{cases}a=f^{\prime}(c,E,k),&n\geq N\\ e=\text{Search}(q),&n<N\end{cases}$
@@ 4.1 Datasets
In our study, we utilized four datasets from prior research that align with our experimental setup: FacTool (Chern et al., 2023), FELM (Chen et al., 2023), Factcheck-Bench (Wang et al., 2023), and BingCheck (Li et al., 2024b).
FacTool and FELM provide factuality claims across multiple domains. From these, we selected instances requiring world knowledge for verification, which we refer to as FacTool-QA and FELM-WK, both with binary annotations (True or False). Factcheck-Bench and BingCheck, however, pose a four-label classification task: supported, partially supported, not supported, and refuted. Following Iqbal et al. (2024), we merged supported and partially supported into True, treated refuted as False, and excluded not supported. Given BingCheck’s imbalance (3,581 True, 42 False), we sampled 100 True claims for our test set. For FELM-WK, we used un-split claims to avoid loss of context. Dataset statistics are in Table 2.
@@ 4.3 Compared Fact-checking Frameworks
We select several SOTA fact-checking frameworks for comparison. Additionally, we introduce two baseline models: Random and Always True/False. To further assess the impact of LLM reasoning in fact-checking, we include an ablation setting.
@@
Table 2: Statistics of the datasets after processing.
@@ 5.1 Preliminary studies
Framework LLM LLM Cost (\$) Search Cost (\$) Time (hrs)
@@
Table 7: Time and USD cost for evaluating the total 559 atomic claims in FacTool-QA, FELM-WK, and BingCheck. We use SerperAPI for FacTool, Safe and Fire for search, while Factcheck-GPT has its own implemented scrapping technique.
@@ 5.2 Comparisons to Other Frameworks
We present a performance comparison of our framework against other frameworks in Table 6 and a cost analysis in Table 7. As shown, all frameworks exhibit similar performance, with a small gap of approximately 0.2 to 0.5. Our framework, using GPT-4o, performs slightly better, achieving superior results on 7 out of 18 metrics, followed closely by Safe with GPT-4o at 6 metrics. This suggests that all frameworks can effectively perform fact-checking for most claims, although they may encounter difficulties with challenging examples, which we analyze further in § 6. In terms of model size, GPT-4o generally outperforms GPT-4o-mini across most frameworks, indicating that larger models are more effective in detecting misinformation. However, the performance improvement is modest, as it comes with an average increase of 16.7 times in LLM costs and an additional 3-fold increase in search costs when using Fire. When considering all frameworks with GPT-4o-mini, Fire demonstrates cost savings of 7.6 times in LLM expenses and 16.5 times in search costs compared to other frameworks. Therefore, we argue that Fire, when paired with GPT-4o-mini, presents a compelling option for the large-scale production of fact-checking systems.
@@
Figure 2 illustrates the impact of reasoning on the number of web searches conducted by GPT-4o and GPT-4o-mini when using BingCheck. Notably, GPT-4o-mini demonstrates a high level of confidence in making verifications when it is allowed to articulate its reasoning process, resulting in the majority of judgments being made without any searches. Conversely, when not permitted to express its reasoning, there is a significant decrease in the number of instances with zero searches; most cases now involve at least one search, indicating a marked reduction in GPT-4o-mini’s confidence in its judgments. This observation aligns with previous findings that the presence of CoT reasoning correlates with increased confidence in the model’s answers (Wang and Zhou, 2024). While GPT-4o also shows a decline in confidence when it is not allowed to search, the decrease is less pronounced than that observed in GPT-4o-mini. By combining the performance and cost results presented in Table 6 and Table 7, we find that, in the absence of a reasoning process, the costs associated with LLMs can be reduced through fewer completion tokens. However, this reduction leads to increased search costs, resulting in overall performance that is inferior to scenarios in which the models are permitted to engage in step-by-step reasoning. Furthermore, the step-by-step reasoning approach facilitates more effective error analysis.
@@
Figure 2: The effect of reasoning on the number of searches using GPT-4o and GPT-4o-mini within Fire on BingCheck. The x-axis shows the number of web searches, while the y-axis denotes the number of instances.
@@ 7 Conclusions and Future Work
We identify several promising directions for future work: (1) Integrating memory banks to store verification results, enabling the system to leverage previous results instead of repeatedly executing the entire process; (2) Expanding the system to support additional modalities, such as code and images.
@@
We acknowledge certain limitations in this work that we plan to address in future research. First, to maintain the efficiency of our framework, we implement the Final Answer or Next Search Query in a compact manner to retrieve evidence, assess confidence on knowledge, and verify the final answer within a single step. Ideally, this can be implemented via separate process which involves a standalone confidence estimation step. This would offer greater flexibility and interpretability. We plan to leave this to our future work. Second, for fair comparison across multiple fact-checking datsets, our system adopts a binary labeling scheme (True or False) and standarlizes the labels across datasets. However, this may not fully reflect the complexity of factual labels in the real world. We intend to incorporate fine-grained labeling schemes in the future.

### v2

@@ 1 Introduction
Large language models (LLMs) have demonstrated exceptional performance across a wide range of tasks, including both language comprehension and generation (Zhao et al., 2023; Xie et al., 2023a). Consequently, LLMs are now widely applied in various domains (Xie et al., 2023b), and many users increasingly rely on the information they provide. However, this reliance is problematic, as LLMs are capable of producing outputs that are highly confident but factually incorrect, highlighting the critical need for robust fact-checking systems (Akhtar et al., 2023). However, fact-checking the entire output of LLMs in a single step is highly challenging. To address this, Min et al. (2023) proposed decomposing the content into multiple atomic claims, each of which can be individually verified. While this approach simplifies the fact-checking process, assessing the veracity of these atomic claims remains complex, especially when many require sourcing evidence from the web. Indeed, identifying the most relevant evidence online is a key challenge in fact-checking pipelines (Wang et al., 2024a).
To address this issue, conventional methods, such as FacTool and Factcheck-GPT (Chern et al., 2023; Wang et al., 2024a), frame the problem as a question-answering task, as illustrated on the left side of Figure 1. In these approaches, an LLM is prompted to generate N relevant questions, which are then used as search queries by a web search tool. The search results serve as evidence for LLM to determine the factuality of the claim. However, we argue that this process is inefficient in two key aspects. First, it underutilizes the internal knowledge already embedded in LLMs during pre-training. For claims involving common knowledge or widely known events, the LLM could confidently assess the claim without relying on external information. Second, generating multiple search queries concurrently does not align with the typical human reasoning process during search (Hu et al., 2023). Humans tend to begin with an initial query, gather information, and then refine their perspective on the claim, which often leads to the formulation of more effective follow-up queries.
@@ 2 Related Work
Despite the remarkable capabilities of LLMs (Brown et al., 2020; OpenAI, 2023; Zhao et al., 2023), the auto-regressive learning objective does not inherently offer strong guarantee or enforce the learning of factual accuracy in the training process, making these models produce content that deviates from real-world facts (Wang et al., 2024b). On average, there are 5%-10% false claims in responses of GPT-4 (OpenAI, 2023) and LLaMA-2 (Touvron et al., 2023) on world-knowledge questions (Iqbal et al., 2024). Retrieval-augmented generation (Guu et al., 2020) and post-generation fact-checking are essential for ensuring accurate knowledge dissemination. Retrieving highly relevant information plays a pivotal role in both guiding generation as a reference and determining verification results in fact-checking systems (Wang et al., 2024a).
@@
The recent advancements in LLMs have spurred significant research on LLM-powered agents, which are capable of reasoning about their environment and making decisions by either invoking external tools or performing internal actions (Wang et al., 2024c). These agent frameworks typically consist of several components, including reasoning, tool usage, memory, and multi-agent debate (Masterman et al., 2024), many of which can be seamlessly integrated into fact-checking pipelines to enhance the performance of traditional fact-checking systems. For example, recent works have endowed systems with the ability to call external tools, such as search engines (Chern et al., 2023; Wang et al., 2024a; Wei et al., 2024; Cheng et al., 2024), recognizing that many claims in the field require additional information for verification. During the verification stage, Sun et al. (2024) proposed a Markov Chain-based multi-agent debate approach to ensure more rigorous verification by enabling collaborative decision-making among agents based on retrieved evidence. Our work differs from previous approaches by combining the evidence retrieval and verification stages, leveraging agents’ reasoning and tool-use capabilities to more closely simulate human cognitive processes in fact-checking.
@@ 3 Framework
Assessing the factual accuracy of long-form text presents significant challenges (Min et al., 2023). To address this, prior approaches have broken down the text into individual checkworthy claims (Chern et al., 2023). These sentences, referred to as atomic claims, are fact-checked individually, with their factuality scores aggregated to evaluate the overall factual accuracy of the original text. Previous research indicates that verifying the factuality of atomic claims is the most challenging step in this process (Wang et al., 2024a). Our work therefore focuses on this critical task: determining the factual accuracy of individual atomic claims, classifying each as either True or False.
@@ 3.1 Fire
We introduce a unified method, Final Answer or Next Search Query $f(\cdot)$, which integrates claim verification with search query generation. Given an atomic claim $c$, this component decides whether to produce a final answer $a$ or generate an additional search query $q$. This decision is guided by both an external evidence set $E$, derived from search results, and the internal knowledge $k$ of the language model, acquired during pre-training. At the outset, no evidence has been retrieved, meaning that the evidence set $E$ is initially empty. Consequently, the decision relies solely on the internal knowledge $k$. As shown in Equation 1, we incorporate confidence estimation into the reasoning process to determine the next action. If the model’s confidence is sufficiently high, it outputs a final answer $a$; otherwise, it generates an additional query $q$.
$f(c,E,k)=\begin{cases}a,&\text{ if confident}\\ q,&\text{ if not confident}\end{cases}$ (1)
@@
When the language model determines that a web search is necessary and issues a search query $q$, we retrieve results using Google Search via the SerpAPI [^1] https://serpapi.com, following prior work (Wei et al., 2024). This API returns the retrieved snippets as a single string, which we use as the new evidence $e$. We then append $e$ to the existing evidence set $E$ to form the updated evidence set $E^{\prime}$ for the next iteration, as shown in Equation 2.
$E^{\prime}=E\cup e,e=\text{Search}(q)$ (2)
@@
Due to the inherent difficulty of confidently verifying certain claims, even with supplementary evidence, we impose an upper limit on the number of retrieval steps. As shown in Equation 3, once this limit is reached, the model performs a final verification $f^{\prime}(\cdot)$ based on all previously retrieved evidence. The detailed prompt for this process is provided in Appendix B.
$\begin{cases}a=f^{\prime}(c,E,k),&n\geq N\\ e=\text{Search}(q),&n<N\end{cases}$ (3)
@@ 4.1 Datasets
In our study, we utilized four datasets from prior research that align with our experimental setup: FacTool (Chern et al., 2023), FELM (Chen et al., 2023), Factcheck-Bench (Wang et al., 2024a), and BingCheck (Li et al., 2024b). FacTool and FELM provide factuality claims across multiple domains. From these, we selected instances requiring world knowledge for verification, which we refer to as FacTool-QA and FELM-WK, both annotated with binary labels (True or False). Our selection was motivated by the need to focus on claims that challenge models to use external knowledge, a critical aspect of factual verification.
For Factcheck-Bench and BingCheck, we consolidated the original four-label classification (supported, partially supported, not supported, refuted) into a binary format by merging supported and partially supported into True, treating refuted as False, and excluding not supported. This binarization aligns these datasets with the others and simplifies evaluation, focusing on clear-cut factuality decisions. We sampled a subset of BingCheck due to its class imbalance (3,581 True claims versus 42 False claims), selecting 100 True claims for our test set. This sampling was essential to create a more balanced and manageable test set, ensuring that evaluation metrics reflect performance on both classes without being dominated by the majority class. In FELM-WK, we retained un-split claims to maintain contextual integrity, which is crucial for accurate verification. Full dataset statistics are provided in Table 1.
Table 1: Statistics of the datasets after processing.
@@ 4.3 Compared Fact-checking Frameworks
We select several SOTA fact-checking frameworks for comparison. Additionally, we introduce two baseline models: Random and Always True/False. To further assess the impact of LLM reasoning and evidence retrieval in fact-checking, we include two ablation settings: Fire (No Reason) and Fire (No Search).
@@
Fire (No Search)
employs the same framework as Fire; however, it is not permitted to invoke the search tool. This configuration is designed to evaluate the model’s ability to perform fact-checking without retrieving any supporting evidence.
@@ 5.1 Preliminary studies
Framework LLM LLM Search Time
@@
Fire (No Search) GPT-4o 1.70 - 1.03
GPT-4o-mini 0.11 - 1.34
Table 7: LLM/Search cost (USD) and time (hrs) for evaluating the total 559 atomic claims in FacTool-QA, FELM-WK, and BingCheck. We use SerperAPI for FacTool, Safe and Fire for search, while Factcheck-GPT has its own implemented scrapping technique.
@@ 5.2 Comparisons to Other Frameworks
We present a performance comparison of our framework against other frameworks in Table 6 and a cost analysis in Table 7. As shown, all frameworks exhibit similar performance, with a small gap of approximately 0.2. Our framework, using GPT-4o, performs slightly better, achieving superior results on 7 out of 18 metrics, followed closely by Safe with GPT-4o at 6 metrics. This suggests that all frameworks can effectively perform fact-checking for most claims, although they may encounter difficulties with challenging examples, which we analyze further in § 6. Regarding the necessity of evidence retrieval in fact-checking, we observe a relatively larger performance drop in FacTool when evidence search is omitted, compared to a smaller drop in FELM-WK and BingCheck. This suggests that FacTool-QA comprises more rare knowledge than GPT-4o-mini, whereas FELM-WK and BingCheck may rely predominantly on common knowledge, for which evidence retrieval is less impactful. Overall, both GPT-4o and GPT-4o-mini perform reasonably well on popular public datasets, highlighting the need for datasets that incorporate more complex claims. In terms of model size, GPT-4o generally outperforms GPT-4o-mini across most frameworks, indicating that larger models are more effective in detecting misinformation. However, the performance improvement is limited, and the associated costs result in an average increase of 16.7 times in LLM expenses and a three-fold increase in search costs when using Fire. Therefore, we argue that cheaper models, such as GPT-4o-mini, are a viable option for performing fact-checking tasks. Furthermore, when considering all frameworks with GPT-4o-mini, Fire achieves additional cost savings, reducing LLM expenses by 7.6 times and search costs by 16.5 times compared to other frameworks. Thus, we contend that Fire, when paired with GPT-4o-mini, offers a compelling solution for the large-scale deployment of fact-checking systems.
@@
Figure 2 illustrates the impact of reasoning on the number of web searches conducted by GPT-4o and GPT-4o-mini tested on BingCheck. Notably, GPT-4o-mini demonstrates a high level of confidence in making verifications when it is allowed to articulate its reasoning process, resulting in the majority of judgments being made without any searches. Conversely, when not permitted to express its reasoning, there is a significant decrease in the number of instances with zero searches; most cases now involve at least one search, indicating a marked reduction in GPT-4o-mini’s confidence in its judgments. This observation aligns with previous findings that the presence of CoT reasoning correlates with increased confidence in the model’s answers (Wang and Zhou, 2024). While GPT-4o also shows a decline in confidence when it is not allowed to search, the decrease is less pronounced than that observed in GPT-4o-mini.
By combining the performance and cost results presented in Table 6 and Table 7, we find that, in the absence of a reasoning process, the costs associated with LLMs can be reduced through fewer completion tokens. However, this reduction leads to increased search costs, resulting in overall performance that is inferior to scenarios in which the models are permitted to engage in step-by-step reasoning. Furthermore, the step-by-step reasoning approach facilitates more effective error analysis.
@@
Figure 2: The effect of reasoning on the number of searches using GPT-4o and GPT-4o-mini within Fire on BingCheck. The shaded area indicates the number of misclassified cases. The x-axis shows the number of web searches, while the y-axis denotes the number of instances.
@@ 7 Conclusions and Future Work
We identify several promising directions for future work, which include: (1) Integrating memory banks to store verification results, allowing the system to reuse previous results instead of repeatedly executing the entire process; (2) Expanding the system to support additional modalities, such as code and images; and (3) Revisiting existing public fact-checking datasets, incorporating personal opinions when addressing ambiguous cases, and adding claims that require rarer and more complex knowledge, where evidence retrieval is essential.
@@
We acknowledge several limitations in this work that we plan to address in future research. First, to maintain the efficiency of our framework, we implement the “Final Answer or Next Search Query” mechanism in a compact manner, allowing it to retrieve evidence, assess confidence in knowledge, and verify the final answer within a single step. Ideally, this process could be separated to include a standalone confidence estimation step, which would enhance both flexibility and interpretability. We leave this exploration to future work. Second, to ensure a fair comparison across multiple fact-checking datasets, our system adopts a binary labeling scheme (“True” or “False”) and standardizes labels across datasets. However, this approach may not fully capture the complexity of factual labels in real-world settings. We intend to incorporate fine-grained labeling schemes in future research. Finally, in this study, we rely on SerpAPI with its default settings. While we did not investigate in detail how evidence is retrieved, we believe future work could explore this aspect further to optimize the selection of the most relevant evidence for a given claim.
@@ Acknowledgments
We thank our reviewers for their valuable reviews and feedback, which significantly contributed to the improvement of our paper.
@@ 5. Format your final answer as a JSON object in the following structure:
Appendix C Effect of Reasoning
We additionally include figures to illustrate the effect of reasoning on two other datasets: FacTool-QA (Figure 3) and FELM-WK (Figure 4), supplementing Figure 2. These figures demonstrate that both GPT-4o and GPT-4o-mini are influenced by explicitly stating their reasoning process, with GPT-4o-mini showing a consistent impact across all datasets, not just BingCheck. Furthermore, when comparing these datasets, we observe that the models appear most confident on FELM-WK compared to the other two datasets. As a result, even in the absence of explicit reasoning, they do not perform any searches to verify the claims.